OpenStudy (anonymous):

need HELP! Use the van der Waals equation of state to calculate the pressure of 2.10 mol of NH3 at 485 K in a 5.50-L vessel. Use the ideal gas equation to calculate the pressure under the same conditions.

OpenStudy (fretje):

ideal gas equation pV=nRT p pressure in SI (international system) or thus Pascal V volume in m³ (cubic meter) n amount of moles R gas constant 8.31441 T absolute temperature

OpenStudy (fretje):

i get thus for the approximation with ideal gas law 1539677.561 Pascal

OpenStudy (fretje):

rest i will try tomorrow to solve.

OpenStudy (matt101):

The van der Waals equation of state is similar to the ideal gas law but it includes the corrections for pressure and volume which are used when not dealing with a real, rather than an ideal, gas. The van der Waals equation of state is the following: \[\left[ P+a \left(n \over V \right)^2 \right] \left({V \over n}-b \right)=RT\] Where a and b are correctional factors that depend on the gas being studied. You will probably have a table or something that will tell you the values for a and b for NH3. The ideal gas equation is what @fretje mentioned above.

OpenStudy (fretje):

@matt101 I get from http://en.wikipedia.org/wiki/Van_der_Waals_equation multiple forms of the state equation, with between them different meaning of the parameters, one should be careful not to mistake, no?: in one form there is a and b, in another a and b are \[ a = N _{A}^{2}.a ^{'}\] and \[ b = N _{A}^{2}.b ^{'}\] .

OpenStudy (matt101):

@fretje The form of the equation I wrote above is the most common one that should be used in this situation. The values of a and b listed in tables are almost always for the equation in that form, so there is probably no need to make a conversion using a' and b'.

OpenStudy (matt101):

I should add that what I wrote above is the same as the second equation on the wikipedia page, I just have my equation divided by n on both sides.